Ascertain that (del)(r^n)=nr^n-2 vector r
1
Expert's answer
2019-10-28T10:44:05-0400

For a function "f(r)" , which depends solely on absolute value of radius-vector "\bold r", the formula for the gradient is the following: "\nabla f(r) = \frac{\partial f(r)}{\partial r} \frac{\bold r}{r}" (this can be shown via direct differentiation w.r.t components or "\bold r"). Hence, "\nabla (r^n) = n r^{n-1} \cdot \frac{\bold r}{r} = n r^{n-2} \bold {r}".